The profit P, in dollars, gained by selling x computers is modeled by the equation P=-5x^2+1000x+5000 How many computers must be sold to obtain a profit of $55,000.00?
Answer by Boreal(15235)

55000=-5x^2+1000x+5000
5x^2-1000x+50000=0
x^2-200x+10000=0
(x-100)^2=0
x=100 computers